# RAP: Retrieval-Augmented Planner for Adaptive Procedure Planning in Instructional Videos

### v1

@@ Abstract
Procedure Planning in instructional videos entails generating a sequence of action steps based on visual observations of the initial and target states. Despite the rapid progress in this task, there remain several critical challenges to be solved: (1) Adaptive procedures: Prior works hold an unrealistic assumption that the number of action steps is known and fixed, leading to non-generalizable models in real-world scenarios where the sequence length varies. (2) Temporal relation: Understanding the step temporal relation knowledge is essential in producing reasonable and executable plans. (3) Annotation cost: Annotating instructional videos with step-level labels (i.e., timestamp) or sequence-level labels (i.e., action category) is demanding and labor-intensive, limiting its generalizability to large-scale datasets.In this work, we propose a new and practical setting, called adaptive procedure planning in instructional videos, where the procedure length is not fixed or pre-determined. To address these challenges we introduce Retrieval-Augmented Planner (RAP) model. Specifically, for adaptive procedures, RAP adaptively determines the conclusion of actions using an auto-regressive model architecture. For temporal relation, RAP establishes an external memory module to explicitly retrieve the most relevant state-action pairs from the training videos and revises the generated procedures. To tackle high annotation cost, RAP utilizes a weakly-supervised learning manner to expand the training dataset to other task-relevant, unannotated videos by generating pseudo labels for action steps. Experiments on CrossTask and COIN benchmarks show the superiority of RAP over traditional fixed-length models, establishing it as a strong baseline solution for adaptive procedure planning.
@@ 1 Introduction
Figure 1: Examples of adaptive procedures in instructional videos. The length of procedure naturally varies according to the differences between initial and goal states, even though the states depict the same task. However, previous works hold an unrealistic assumption that the sequence length is fixed and known in advance.
Procedure planning in instructional videos aims to generate an executive series of actions for completing tasks, transitioning from initial state to final visual state [4, 2, 19, 24, 22, 23, 12, 16]. While humans excel at such tasks, machine learning agents face challenges in achieving similar proficiency. Success in procedure planning requires an agent’s deep understanding of actions, procedure states, involved objects, and their temporal relationships. Narrowing this performance gap is key to advancing autonomous systems, equipping them to adeptly perform complex, instruction-based tasks like cooking or repair.
@@ 2.2 Sequence Generation with Auto-regressive Transformers
Recent advancements in sequence generation using Transformer-based models [21, 17] have established remarkable results in tasks involving sequence modeling, e.g. natural language modeling [17]. Procedure planning in its core, is a goal-conditioned sequence generation problem, thereby dirrectly benefiting from the state-of-the-art sequence generation models. Leveraging these developments, our work incorporates an auto-regressive Transformer [17] into the RAP architecture.
@@ 3.2 Model: Retrieval-Augmented Planner
Figure 2: Our framework overview: Visual observations $o_{s},o_{g}$ and action steps $\{a_{t}\}_{t=1}^{T}$ are encoded to $v_{s},v_{g}$ and $\{l_{t}\}_{t=1}^{T}$. A Task-Classifier takes $o_{s}$ and $o_{g}$ to predict a task class and its corresponding representation $\hat{c}$ which in turn alongside $v_{s}$, $v_{g}$, and a learnable START token $l_{\texttt{START}}$, is fed into the auto-regressive base planner $F_{base}$ for sequential action embedding prediction, using positional embeddings $\{p_{t}\}_{t=1}^{T}$. The contrastive loss $\mathcal{L}_{\text{base}}$ is calculated by comparing predictions $\{\hat{l}_{t}\}_{t=1}^{T+1}$ to $\{l_{t}\}_{t=1}^{T+1}$. For each prediction at position $t$, the retrieval component, based on the context $[v_{s},v_{g},\hat{c},\hat{l}_{t}+q_{t}]$ (with $q_{t}$ being a positional embedding at $t$) retrieves the closest K key-value pairs for kNN linear interpolation of the probability distribution $P_{\text{all}}$, deriving at the cross-entropy loss $\mathcal{L}_{\text{all}}$.
@@
Retrieval Planner. To improve long-range sequence modeling, our model integrates a direct retrieval mechanism based on a training datastore, utilizing its state-action temporal dependency information, explicitly. This enhancement is achieved by augmenting the base planner in a manner similar to [10, 7], by adding a retrieval-augmented memory module, which learns the state-action pairs in the training data as memory, by modeling states, as a context for the actions.
@@ 4 Experiments
We experiment on two datasets and employ four metrics to validate the efficacy of our proposed model.
@@ 4.4 Comparison with State-of-The-Art Baselines
Table 1: Assessment of procedure planning on CrossTask for action-horizons of $T=3$ and $T=4$. We denote the vision supervision by V, the language supervision by L, and the addition of weakly-supervised training by +wL.
@@
RAP (ours with unannotated data) L 30.51 62.68 75.92 17.10 55.91 71.33
@@
Our study conducts a comprehensive comparison of our new framework, RAP, against established models in procedure planning, using CrossTask and COIN datasets. We refer to our framework without the the retrieval-based memory augmentation, i.e., base planner, as BP (as opposed to RAP). We categorize previous approaches based on their supervision: language (L) or vision (V). Our proposed model’s performance compared to the previous works, is detailed in Tab. 1 for CrossTask and Tab. 2 for COIN.
@@
When compared to language-supervised models like P3IV, E3P, and Skip-Plan, RAP demonstrates superior performance on CrossTask. For T=3, RAP exhibits improvements of 5.99%, 2.83%, and 0.38% respectively, and for T=4, it shows enhancements of 3.56%, 0.47%, and 1.4% in SR. Notably, PDPP [23], a recent model with superior performance to RAP, has been excluded from this comparison due to its distinct methodology. PDPP utilizes a 3-second window following the onset and preceding the offset of action steps to capture visual states, which we believe gives it an upper hand in accessing action-related information, compared to the previous models. Conventionally initial and final visual states are captured using a 1-second window centered around the start and end time of actions. In the supplementary material, we provide RAP’s performance compared to PDPP when incorporating a similar methodology for defining visual states, demonstrating RAP’s superiority over PDPP. We observe a similar superiority in performance, when evaluating the model on the COIN dataset as shown in Tab. 2.
@@ 0.C.3 Ablation of the Retrieval Planner Architecture
0.C.4 Investigating the Impact of Visual State Definition
In our implementation of RAP, we have adopted a methodology akin to previous studies [4, 24, 22, 12] for capturing visual states. We define initial and final visual states using a 1-second window centered around action start and end times, which differs from PDPP’s [23] approach. To assess the influence of visual state window duration, we conducted experiments similar to PDPP’s methodology. PDPP utilizes a 3-second window following action step onset and preceding its offset, offering an advantage in accessing action-related information compared to earlier models. Table 12 presents RAP’s performance relative to PDPP when employing a similar methodology for defining visual states, demonstrating RAP’s superiority.
Table 12: Comparison of RAP’s performance with PDPP’s using a similar approach for defining visual states, on the CrossTask dataset with T=3.
Model SR (%) mAcc (%) mIoU (%)
PDPP 37.20 64.67 66.57
RAP (w/o unannotated) 38.51 65.18 77.44

### v2

@@ Abstract
Procedure Planning in instructional videos entails generating a sequence of action steps based on visual observations of the initial and target states. Despite the rapid progress in this task, there remain several critical challenges to be solved: (1) Adaptive procedures: Prior works hold an unrealistic assumption that the number of action steps is known and fixed, leading to non-generalizable models in real-world scenarios where the sequence length varies. (2) Temporal relation: Understanding the step temporal relation knowledge is essential in producing reasonable and executable plans. (3) Annotation cost: Annotating instructional videos with step-level labels (i.e., timestamp) or sequence-level labels (i.e., action category) is demanding and labor-intensive, limiting its generalizability to large-scale datasets. In this work, we propose a new and practical setting, called adaptive procedure planning in instructional videos, where the procedure length is not fixed or pre-determined. To address these challenges, we introduce Retrieval-Augmented Planner (RAP) model. Specifically, for adaptive procedures, RAP adaptively determines the conclusion of actions using an auto-regressive model architecture. For temporal relation, RAP establishes an external memory module to explicitly retrieve the most relevant state-action pairs from the training videos and revises the generated procedures. To tackle high annotation cost, RAP utilizes a weakly-supervised learning manner to expand the training dataset to other task-relevant, unannotated videos by generating pseudo labels for action steps. Experiments on CrossTask and COIN benchmarks show the superiority of RAP over traditional fixed-length models, establishing it as a strong baseline solution for adaptive procedure planning.
@@ 1 Introduction
Figure 1: Examples of adaptive procedures in instructional videos. Procedure length varies with differences between initial and goal visual states, even for the same task. Previous works assume a fixed, known sequence length, requiring a separate model for each action-horizon. This figure contrasts traditional procedure planning (A) with our adaptive approach (B), which trains a unified model without knowing the procedure length in advance.
Procedure planning in instructional videos aims to generate an executive series of actions for completing tasks, transitioning from initial visual state to final visual state [4, 2, 19, 24, 22, 23, 12, 16]. While humans excel at such tasks, machine learning agents face challenges in achieving similar proficiency. Success in procedure planning requires an agent’s deep understanding of actions, procedure states, involved objects, and their temporal relationships. Narrowing this performance gap is key to advancing autonomous systems, equipping them to adeptly perform complex, instruction-based tasks like cooking or repair.
@@ 2.2 Sequence Generation with Auto-regressive Transformers
Recent advancements in sequence generation using Transformer-based models [21, 17] have established remarkable results in tasks involving sequence modeling, e.g. natural language modeling [17]. Procedure planning in its core, is a goal-conditioned sequence generation problem, thereby directly benefiting from the state-of-the-art sequence generation models. Leveraging these developments, our work incorporates an auto-regressive Transformer [17] into the RAP architecture.
@@ 3.2 Model: Retrieval-Augmented Planner
Figure 2: Framework overview: Visual observations $o_{s}$ and $o_{g}$ are encoded to $v_{s}$ and $v_{g}$. A Task-Classifier predicts a task class and its representation $\hat{c}$ using $v_{s}$ and $v_{g}$. These, along with a learnable START token, are fed into the auto-regressive base planner $F_{base}$ for sequential action embedding prediction. For each prediction at position $t$, the retrieval component uses the predicted context vector $\varsigma_{t}$ to estimate a probability distribution $P_{\text{retr,t}}$ over the closest $K$ key-value pairs in memory for kNN linear interpolation with $P_{\text{base,t}}$ to estimate the final probability distribution $P_{\text{all,t}}$.
Figure 3: The Retrieval-Augmented Planner module uses the predicted context vector $\varsigma_{t}$ to retrieve the $K$ closest keys and their values from memory, estimating a probability distribution $P_{\text{retr,t}}$ over these entries for next action probability estimation $P_{\text{all,t}}$.
@@
Retrieval Planner. To improve long-range sequence modeling, our model integrates a direct retrieval mechanism based on a training datastore, utilizing its state-action temporal dependency information, explicitly. This enhancement is achieved by augmenting the base planner in a manner similar to [10, 7], by adding a retrieval-augmented memory module, which learns the state-action pairs in the training data as memory, by modeling states, as a context for the actions. Fig. 3 illustrated the workings of this component.
@@ 4 Experiments
We experiment on two benchmark datasets and employ four metrics to validate the efficacy of our proposed model.
@@ 4.4 Comparison with State-of-The-Art Baselines
Table 1: Assessment of procedure planning on CrossTask for action-horizons $T=3$ and $T=4$. Vision supervision is denoted by V, language supervision by L, and weakly-supervised training by +wL.
@@
RAP (ours with unannotated data) L+wL 30.51 62.68 75.92 17.10 55.91 71.33
Table 2: Comparison of RAP with PDPP, when we align our state definition with PDPP’s for CrossTask and $T=3$.
Models Supervision $SR\uparrow$ $mAcc\uparrow$ $mIoU\uparrow$
PDPP [23] $\clubsuit$ L 37.20 64.67 66.57
RAP (ours w/o unannotated data) $\clubsuit$ L 38.51 65.18 77.44
@@
Our study conducts a comprehensive comparison of our new framework, RAP, against established models in procedure planning, using CrossTask and COIN datasets. We refer to our framework without the the retrieval-based memory augmentation, i.e., base planner, as BP (as opposed to RAP). We categorize previous approaches based on their supervision: language (L) or vision (V). Our proposed model’s performance compared to the previous works, is detailed in Tabs. 1 and 2 for CrossTask, and Tab. 3 for COIN.
@@
When compared to language-supervised models like P3IV, E3P, and Skip-Plan, RAP demonstrates superior performance on CrossTask. For T=3, RAP exhibits improvements of 5.99%, 2.83%, and 0.38% respectively, and for T=4, it shows enhancements of 3.56%, 0.47%, and 1.4% in SR. Tab. 2 compares our model with PDPP [23], which employs a different procedure state definition. For fairness, we evaluate our model under the PDPP setting, marked with $\clubsuit$, showcasing its superior performance. PDPP uses a 3-second window following the onset and preceding the offset of action steps to capture visual states, giving it an advantage in accessing action-related information compared to previous models. Traditionally, a 1-second window centered around the start and end of actions is used. By aligning our state definition with PDPP’s, our RAP model outperforms PDPP on all metrics, as shown in the table. We observe a similar superiority in performance, when evaluating the model on the COIN dataset as shown in Tab. 3.
@@ Appendix 0.A Grounding Algorithm
Fig. 7 shows an example of grounding instructional videos to the action plans, illustrating the generation of pseudo time-stamps for action start and end points.
Figure 7: An example of grounding action steps to video to obtain action start and end states.
@@ Appendix 0.B Implementation Details
We used a single NVIDIA TITAN RTX GPU with $24$G of memory for training and evaluation. The inference time is $0.3$s per sample. The training time is $9.1$ hours and $3.8$ hours for stage 1 and stage 2, respectively.
@@ 0.C.3 Ablation of the Retrieval Planner Architecture
Tab. 10 shows RAP components’ performance. The Retrieval head alone under-performs the Base Planner (BP) but can improve BP when combined as RAP.
@@
@R2 Performance of retrieval head $P_{retr}$. The below table shows RAP components’ performance. The Retrieval head alone underperforms the Base Planner (BP) but can improve BP when combined as RAP.
Table 10: Ablation study on the performance of the retrieval head alone on CrossTask for the setting of variable-length action horizon.
Model SR (%) mAcc (%) mIoU (%) mES (%)
BP $P_{base}$ 18.54 47.85 69.16 62.64
Retrieval head $P_{retr}$ 11.42 23.28 43.67 35.96
RAP (w/o unannotated data) 20.28 49.41 69.95 62.79